Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability question.Mismatch in answer in 2 diff approaches

Expert replies
by vmdb » Thu Jan 21, 2010 6:36 pm
Question: A florist has 2 Azaleas, 3 Buttercups, and 4 Petunias. She put two flowers together at random in a bouquet. However a customer does not want two of the same flower. What is the probability that the florist does not have to change the bouquet ?

I tried answering the above problem in two ways.

1) Prob of having 1 azalea&1 buttercup in a bouquet OR 1 buttercup&1 petunia in a bouque OR 1 Petunia&1 azalea in a bouquet.

=> 2/9*3/8 + 3/9*4/8 + 4/9*2/8 = 26/72


2) 1 - (Prob of having 1 azalea&1azalea OR 1 buttercup&1 buttercup OR 1petunia&1 petunia)

=> 1 - (2/9*1/8 + 3/9*2/8 + 4/9*3/8) = 1 - 20/72 = 13/18



My question is why am I getting TWO different answers. They both have to be the same. 13/18 is the correct answer, as per the Guide. What is the mistake I 'm doing in my FIRST approach
Join the discussion
Source: — Problem Solving |

by ace_gre » Thu Jan 21, 2010 7:43 pm
Hi , Here is my take on the first approach..

2/9*3/8 *2+ 3/9*4/8 *2+ 4/9*2/8 *2

*2 because there are two ways of selecting one azalaa and one buttercup==>
1 azalea & 1 buttercup
1 buttercup & 1 azalea

Same for the other combinations. So 26/72 * 2 = 13/18.
Not an issue in second approach because both the flowers are of the same type...

Hope this helps!
Join the discussion

by vmdb » Fri Jan 22, 2010 8:46 am
I didn't really understand, how a) picking first: Azalea & picking second: Buttercup differ from b)picking first: Buttercup & picking second: Azalea. No matter how one picks up, contents of the bouquet remains the same.

I was thinking abt it yesterday, got a different way of analyzing the problem

Prob [(Azalea first AND (Buttercup OR Petunia) second) OR (Buttercup first AND (Azalea OR Petunia) second) OR (Petunia first AND (Buttercup OR Azalea) second)]

That evaluates to

2/9*(3/8 + 4/8) + 3/9*(2/8 + 4/8) + 4/9*(2/8 + 3/8) = 13/18

Anything wrong in this approach.
Join the discussion